A chemistry graduate student is given 250.mL of a 0.60M ammonia NH3 solution. Ammonia is a weak base with =Kb×1.810−5. What mass of NH4Br should the student dissolve in the NH3 solution to turn it into a buffer with pH =9.35?
Answer #1

Kb = 1.8*10^-5

pKb = - log (Kb)

= - log(1.8*10^-5)

= 4.745

POH = 14 - pH

= 14 - 9.35

= 4.65

use formula for buffer

pOH = pKb + log ([NH4Br]/[NH3])

4.65 = 4.7447 + log ([NH4Br]/[NH3])

log ([NH4Br]/[NH3]) = -0.0947

[NH4Br]/[NH3] = 0.804

[NH4Br]/0.6 = 0.804

[NH4Br] = 0.4824

volume , V = 2.5*10^2 mL

= 0.25 L

use:

number of mol,

n = Molarity * Volume

= 0.4824*0.25

= 0.1206 mol

Molar mass of NH4Br,

MM = 1*MM(N) + 4*MM(H) + 1*MM(Br)

= 1*14.01 + 4*1.008 + 1*79.9

= 97.942 g/mol

use:

mass of NH4Br,

m = number of mol * molar mass

= 0.1206 mol * 97.94 g/mol

= 11.81 g

Answer: 11.8 g